5m long beam P=90 kn Find the deflection at point p using Virtual Work Method

5m long beam

P=90 kn

Find the deflection at point p using Virtual Work Method
Answer #1

MA-PEL XXL P-RA Reaction Ey=o. PA=P , PATI Empo - MA + PXL O - MAX 1XL zo - MA=PL (G) - MAEL LILL Ma- P*X-PL mx = x-L NOLO 4P-90KN Лу - - Ру mo- me = -1 = 900 By virtual work I Mmda ox) (-3) da A Immax - (- 902) (3) de 90823 - 90x23 BEL 4 +302 km. iof virtual work unit load Principal FOR Beam & frame Now Foom Flexual Fermula ņķ do = 1 de From Principle of virtual work. 1.